# Restoring Accessibility During Urban Rail Disruptions via Bus Network Redesign

### v1

@@ 1 Introduction
Urban Rail Transit (URT) plays a pivotal role in metropolitan areas by alleviating suburban-to-downtown congestion, offering high energy efficiency, and convenient service across various weather conditions [1, 2]. However, URT systems face operational challenges, to maintain high reliability [3]. Notable incidents, such as the July 2017 fire at New York City’s 145th Street Station caused by track debris, have led to significant service disruptions for more than 2 hours during morning peak  [4]. Even the highly efficient Hong Kong Mass Transportation Railway (MTR), boasting a 99.9% on-time performance, encounters approximately 250 disruptions annually [5]. In Paris, the RER B line typically serves nearly 200,000 passengers on a weekday. However, for maintenance work, long segments of the line have been closed in recent years, over multiple hours. Replacement buses operating to remediate this closure have proven to be insufficient, forcing many commuters to seek alternative routes or reduce their travel [6].
@@
In city centers with dense public transport (PT) networks, passengers are typically able to find alternative lines. With sparse public transport, nonetheless, the effect of URT disruption will be much worse. Due to the lack of alternatives, passengers in these areas might be left with no mobility options and have severe difficulties in performing their activities. To take these aspects into account, the usual level of service metrics (average waiting times, travel times, etc.) are not sufficient. What is more relevant to passengers is instead accessibility. Stated in simple terms, accessibility indicators measure how many opportunities per hour can be reached from a certain location [7]. Opportunities can be workplaces, schools, healthcare facilities or any other location that must be reached to perform activities relevant for the users. Nonetheless, remediation strategies applied during disruptions do not generally take accessibility into account. Indeed, conventional remediation strategies consist of providing substitute buses and primarily focus on minimizing operational metrics while overlooking the crucial factor of area accessibility.
@@
To fill these gaps, we propose a bus network redesign method to restore accessibility during URT disruptions. In order to serve the impacted passengers, our approach decides which lines to extend, calculates the routes of these extensions, and re-optimizes the allocation of buses across lines. We show (and explain) the superiority of our approach with respect to usual practice, which consists in just replacing the interrupted URT line with a “replacement bus”, which just stops at the original disrupted stops. Due to the problem’s complexity, we further develop a two-stage heuristic optimization method that generates solutions within seconds. We validate our approach using two case studies of rail disruptions in two French towns in the suburbs of the Paris region. Results demonstrate that our method recovers the accessibility loss during disruptions much more efficiently than conventional replacement buses, even without adding buses, just by appropriately extending existing lines and reallocating the existing fleet of buses.
@@ 2 Related Work
Due to its inherent complexity, Urban Rail Transit (URT) disruption management has been largely studied in the literature. Shalaby et al. [10] reviewed recovery models and algorithms for real-time railway disruption management, which involves timetable adjustment, substitute bus scheduling, and crew rescheduling. The in-depth study on substitute bus scheduling was done by Zhang et al. [11], their model incorporated the additional costs of mobilizing the substitute bus fleet into the objective function, offering a more comprehensive approach to optimizing services.
@@ 2.2 Work Related to Replacement Buses
An approach commonly employed in practice [6], to remediate the disruption of a rail line, is to employ extra-buses in order to travel across the stops of the disrupted rail [20, 21].
Previous literature mainly focused on minimizing transit operating time, passenger discomfort, and system cost. In the same sense, Wang et al. [13] created a model for bus routes optimization for overall time driving minimization, whereas Cadarso et al. [22] designed an optimization model of bus bridging schedules and vehicle rescheduling from a comprehensive point of view, considering minimization of time of recovery, passenger discomfort, and system costs. Luo and Xu [23] incorporated the stochastic nature of passenger demand and existing rail and bus routes’ backup capacities into a stochastic programming model to minimize expected unmet passenger requirements. To cope with uncertain and heterogeneous bus traveling times, Liang et al. [24] developed a passenger flow and operational cost optimization model.
@@ 2.3 Our Contribution
Despite the amount and the high quality of previous work on Urban Rail Transit disruption remediation, all previous studies have overlooked the most crucial factor of a transportation system, i.e., accessibility [7]. This aspect is particularly important when considering disruptions in suburban regions, where accessibility often tightly depends on one or few rail lines. Due to the scarcity of public transport alternatives, a disruption of a rail line can destroy the accessibility of suburbs, preventing people living or working in those areas from normally performing their activities.
For this reason, to the best of our knowledge, we are the first to propose a bus network redesign method to restore accessibility during Urban Rail Transit (URT) disruptions, particularly suited for suburbs with scarce public transport resources.
@@ 3.1 Tessellation of the Study Area
We partition the study area using a regular tessellation, where the center of each tile is defined as a centroid. Let $\pazocal{C}$ denote the set of all centroids in the study area. A traveler can choose from various transportation modes for a trip between any pair of centroids. For instance, a traveler may walk the entire distance at speed $v_{walking}$, or they could walk from the origin centroid $c_{i}$ to a PT station, then take PT to another station, and finally walk from there to the destination centroid $c_{j}$. Observe that while actual trips may be longer or shorter than centroid-to-centroid distances, these variations tend to balance out across the PT network. This is why we approximate all travel as centroid-to-centroid.
@@ 3.2 Graph Model of PT Network
We model the original PT structure before disruption as graph $\pazocal{G}^{o}=(\pazocal{V}^{o},\pazocal{E}^{o})$. $\pazocal{G}^{o}$ is composed of multiple PT lines. Each line $l$ has a headway $t_{l}$, defined as the distance between two consecutive vehicles expressed in time [25]. The headway can be calculated using the formula $t_{l}=\frac{T_{l}}{N_{l}}$, where $T_{l}$ represents the round trip time along the entire line, and $N_{l}$ denotes the number of vehicles operating on that line (Eq. 4.4 of [25]). A line consists of a sequence of stations connected by edges. The average waiting time for line $l$ is $\frac{t_{l}}{2}$ [26], We define $t(s_{j},s_{j+1})$, the time taken to travel between any two successive stations $s_{j}$ and $s_{j+1}$, and $t_{s_{j}}$, the dwell time at station $s_{j}$. For interchange at station $s_{j}$ from line $l$ to line $l^{\prime}$, we consider an average waiting time $\frac{t_{l^{\prime}}}{2}$ at station $s_{j}$.
@@ 3.3 Problem Statement
In the face of urban rail system disruptions, our research aims to maximize accessibility through a redesign of bus lines. We extend operating bus routes allowing buses of one line to serve both stops of the original line and consolidation points at which passengers concerned by the disruption can be picked up and dropped off. Our optimization problem makes the following five decisions: (a) To which candidate consolidation points should passengers from the disrupted rail stations be suggested to go? (b) To which bus lines should the consolidation nodes be assigned? (c) Which bus line should be extended in order to serve such passengers? (d) When a bus serves multiple consolidation points, what is the optimal order of visits? (e) How to re-assign the fleet of buses among the original and extended lines?
@@
As illustrated in Fig. 2, for any given line $l$, let $t_{l}^{a}$ and $t_{l}^{b}$ represent its terminals. We can activate an extended version of $l$ from terminal $t_{l}^{b}$, which will serve the stops of the regular $l$ as well as consolidation nodes that will be associated with this line. Let’s denote this extended line as $l^{EX,b}$ (red line). Similarly, we can define $l^{EX,a}$ (blue line). Let $\pazocal{L}_{\text{REG}}=\{l_{1},l_{2},...,l_{l}\}$ be the set of regular lines. Let $\pazocal{L}_{\text{EX}}=\{l^{EX,a},l^{EX,b}|l\in\pazocal{L}_{\text{REG}}\}$ the set of all potential extended bus lines.
Table 1: Table of Notation
@@
$n$ Consolidation node index
$d$ Demand point index
$i$, $j$ Indices for nodes
$l$ Bus line index
@@
$\pazocal{N}$ The set of consolidation nodes, i.e., locations where affected passengers who wish to take the redesigned bus can board.
$\pazocal{D}$ Set of disrupted Rail stations; i.e., locations where the affected passengers who wish to take the redesigned bus will be gathered.
$\pazocal{C}$ The set of centroids in the study area.
$\pazocal{L}=\pazocal{L}_{REG}\cup\pazocal{L}_{\text{EX}}$ The set of all bus route lines in the network, where $\pazocal{L}_{\text{REG}}$ means the regular lines and $\pazocal{L}_{\text{EX}}$ includes any extended lines proposed by the optimization model.
@@
$q_{d}$ Passenger demand at disrupted rail station $d$, expressed in number of passengers per unit of time, $\forall d\in\pazocal{D}$.
$cap$ Passenger capacity per bus.
$N^{o}_{l}$ Number of buses operating on original regular line $l$, $\forall l\in\pazocal{L}_{\text{REG}}$.
@@
$D_{max}$ Maximum distance allowed for a disrupted rail station to the suggested consolidation node.
@@
$\pazocal{G}$ The PT structure.
$n_{l}\in\mathbb{Z}^{+}$ Number of additional vehicles per unit of time on line $l$, and its extensions (these add buses can operate on regular line $l$ or extension $l^{EX,a}$ or extension $l^{EX,b}$) serve as a regular line or an extension line, $\forall l\in\pazocal{L}_{\text{REG}}$.
$x_{l}\in\mathbb{Z}^{+}$ Number of vehicles per unit of time on line $l$, $\forall l\in\pazocal{L}$.
$y_{l^{EX}}\in{0,1}$ 1 if extended line $l^{EX}\in\pazocal{L}_{\text{EX}}$ is activated, 0 otherwise.
$w_{nl}\in{0,1}$ 1 if consolidation point $n$ is assigned to the extended line $l$, 0 otherwise, $\forall n\in\pazocal{N},\forall l\in\pazocal{L}_{\text{EX}}$.
$z_{dn}\in{0,1}$ 1 if disrupted rail station $d$ is assigned to consolidation point $n$, 0 otherwise, $\forall n\in\pazocal{N},\forall d\in\pazocal{D}$.
$u_{i}^{l}\in\mathbb{Z}^{+}$ The positional order of node $i$ in line $l$, 0 if $i$ is not visited by $l$, $\forall i\in\pazocal{V},\forall l\in\pazocal{L}$.
$X_{ij}^{l}\in{0,1}$ 1 if there is a direct connection between $i,j$, served by line $l$, 0 otherwise, $\forall i,j\in\pazocal{V},\forall l\in\pazocal{L}$.
@@ 3.4.1 Objective function
where $O_{c_{j}}$ is the number of opportunities within the tile with centroid $c_{j}$ and $T_{\pazocal{G}}(c_{i},c_{j})$ is the shortest travel time between centroids $c_{i}$ and $c_{j}$, considering that trips can combine multiple lines and can also include walk from $c_{i}$ to $c_{j}$ or to/from some intermediate stops; Subscript $\pazocal{G}$ indicates that travel time depends on PT structure $\pazocal{G}$. Indeed, if the PT structure changes (for instance some edges are no more available because of disruption), some shortest paths may change accordingly, resulting in a change in travel time $TG(c_{i},c_{j})$. Let $\pazocal{G}^{o}$ be the original PT structure before disruption, and $\pazocal{G}^{\text{DISR}}$ be the disrupted PT structure. Then, it’s easy to show that:
@@ 3.4.2 Constraints
$\sum_{l\in\pazocal{L}_{\text{EX}}}w_{nl}\leq 1,$ $\forall n\in\pazocal{N}$ (4)
@@
$z_{dn}\leq\sum_{l\in\pazocal{L}_{\text{EX}}}w_{nl},$ $\forall n\in\pazocal{N},\forall d\in\pazocal{D}$ (6)
@@
$\sum_{j\in\pazocal{N}}X_{t_{l}^{a}j}^{l^{EX,a}}=y_{l^{EX,a}}~{};\sum_{j\in\pazocal{N}}X_{t_{l}^{b}j}^{l^{EX,b}}=y_{l^{EX,b}}$ $\forall l\in\pazocal{L}_{\text{REG}}$ (8)
$\sum_{j\in\pazocal{N}}X_{ij}^{l}=\sum_{j\in\pazocal{N}}X_{ji}^{l},$ $\forall i\in\pazocal{N},\forall l\in\pazocal{L}_{\text{EX}}$ (9)
$\sum_{j\in\pazocal{N}}X_{ij}^{l}\leq w_{il},$ $\forall i\in\pazocal{N},\forall l\in\pazocal{L}_{\text{EX}}$ (10)
$\sum_{i\in\pazocal{N}}X_{ij}^{l}\leq w_{jl},$ $\forall j\in\pazocal{N},\forall l\in\pazocal{L}_{\text{EX}}$ (11)
$u_{i}^{l}-u_{j}^{l}+(|N|+2)\cdot X_{ij}^{l}\leq(|N|+1)\cdot w_{il},$ $\forall i,j\in\pazocal{N},i\neq j,\forall l\in\pazocal{L}_{\text{EX}}$ (12)
$2\cdot w_{il}\leq u_{i}^{l}\leq(|N|+1)\cdot w_{il},$ $\forall i\in\pazocal{N},\forall l\in\pazocal{L}_{\text{EX}}$ (13)
$u_{t_{l}^{a}}^{l^{EX,a}}=y_{l^{EX,a}}~{};u_{t_{l}^{b}}^{l^{EX,b}}=y_{l^{EX,b}}$ $\forall l\in\pazocal{L}_{\text{REG}}$ (14)
@@
$x_{l^{EX,a}}\geq y_{l^{EX,a}}~{};x_{l^{EX,b}}\geq y_{l^{EX,b}}$ $\forall l\in\pazocal{L}_{\text{REG}}$ (15)
$x_{l^{EX,a}}\leq M\cdot y_{l^{EX,a}}~{};x_{l^{EX,b}}\leq M\cdot y_{l^{EX,b}}$ $\forall l\in\pazocal{L}_{\text{REG}}$ (16)
$cap\cdot x_{l}\geq\sum_{d\in\pazocal{D}}\sum_{n\in\pazocal{N}}q_{d}\cdot z_{dn}\cdot w_{nl},$ $\forall l\in\pazocal{L}_{\text{EX}}$ (17)
$n_{l}=x_{l}+x_{l^{EX,a}}+x_{l^{EX,b}}-N^{o}_{l},$ $\forall l\in\pazocal{L}_{\text{REG}}$ (18)
$\sum_{l\in\pazocal{L}}n_{l}\leq N_{max}$ (19)
Constraints (4) to (7) handle the allocation relationship between disrupted rail stations and consolidation nodes: Constraint (4) ensures each consolidation node is assigned to at most one extended line; Constraint (5) guarantees each disrupted rail station is allocated to exactly one consolidation node; Constraint (6) ensures disrupted rail stations are only assigned to stops served by extended lines; Constraint (7) limits the distance between a disrupted rail station and its assigned consolidation node to within the maximum allowable distance. Constraints (8) to (14) design the line extensions: Constraint 8 ensures that if an extended line $l$ is activated, for the passing terminal, which is not a consolidation node, there should be exactly one direct connection from $t$ within line $l$. Constraint (9) maintains flow conservation at consolidation nodes and ensures that the number of incoming and outgoing connections is equal. Constraints (10) and (11) limit each node to have at most one outgoing and one incoming edge in each line, respectively, preserving the linear structure of the line. Constraint (12) implements the Miller-Tucker-Zemlin subtour elimination constraint [27], ensuring that the line forms a single path without cycles among consolidation nodes. Lastly, constraint (13) sets the bounds for node position variables, maintaining sequential order, while Constraint (14) specifically designates the starting terminal. Constraints (15) to (19) deal with vehicle allocation and line selection: Constraint (15) and Constraint (16) ensure vehicles are only assigned to a line extension if this extended line is activated; Constraint (17) guarantees the vehicle capacity of each extended line meets demand; Note that $x_{l}$ indicates the number of vehicles operating on the extended line $l$ and $z_{dn}\cdot w_{nl}$ is one if and only if disrupted rail station $d$ is served by extended line $l$; $q_{d}$ can be estimated through analysis of historical passenger flow data and real-time monitoring etc. In our numerical experiments, we employed statistical models; this constraint makes the problem non-linear. Constraint (18) governs the allocation of vehicles across the transit network. For each regular line $l$, it ensures that the number of vehicles on the regular line ($x_{l}$) equals the initial fleet size ($N^{o}_{l}$) plus additional vehicles fleet size ($n_{l}$), minus the number of vehicles on extended lines. Constraint (19) expresses a budget for additional buses to be added. Note that the formulation allows us to reduce the number of buses operating on regular lines in our solution with respect to the PT design before disruption. It may indeed be useful to slightly reduce the service on regular lines to compensate for the loss of accessibility due to disruption.

### v2

@@ 1 Introduction
Urban Rail Transit (URT) plays a pivotal role in metropolitan areas by alleviating suburban-to-downtown congestion, offering high energy efficiency, and convenient service across various weather conditions [1, 2]. However, URT systems face operational challenges to maintain high reliability [3]. Notable incidents, such as the July 2017 fire at New York City’s 145th Street Station caused by track debris, have led to significant service disruptions for more than 2 hours during morning peak [4]. Even the highly efficient Hong Kong Mass Transportation Railway (MTR), boasting a 99.9% on-time performance, encounters approximately 250 disruptions annually [5]. In Paris, the RER B line typically serves nearly 200,000 passengers on a weekday. However, for maintenance work, long segments of the line have been closed in recent years, over multiple hours. Replacement buses operating to remediate this closure have proven to be insufficient, forcing many commuters to seek alternative routes or reduce their travel [6].
@@
In city centers with dense public transport (PT) networks, passengers are typically able to find alternative lines. With sparse public transport, nonetheless, the effect of URT disruption will be much worse. Due to the lack of alternatives, passengers in these areas might be left with no mobility options and have severe difficulties in performing their activities. To take these aspects into account, the usual level of service metrics (average waiting times, travel times, etc.) are not sufficient. What is more relevant to passengers is instead accessibility. Stated in simple terms, accessibility indicators measure how many opportunities per hour can be reached from a certain location [7]. Opportunities can be workplaces, schools, healthcare facilities or any other location that must be reached to perform activities relevant for the users. Nonetheless, remediation strategies applied during disruptions do not generally take accessibility into account. Indeed, conventional remediation strategies consist of providing substitute buses and primarily focus on minimizing operational metrics while overlooking the crucial factor of area accessibility. This represents a significant gap in current practice, as maintaining accessibility is fundamental to preserving residents’ quality of life and social participation during disruptions [8].
@@
To fill this gap, we propose a bus network redesign method to restore accessibility during URT disruptions. Our key innovation lies in making accessibility the primary optimization objective when redesigning bus networks during rail disruptions, rather than focusing solely on conventional operational metrics. In order to serve the impacted passengers, our approach decides which lines to extend, calculates the routes of these extensions, and re-optimizes the allocation of buses across lines. We show (and explain) the superiority of our approach with respect to usual practice, which consists in just replacing the interrupted URT line with a “replacement bus”, which just stops at the original disrupted stops. Due to the problem’s complexity, we further develop a two-stage heuristic optimization method that generates solutions within seconds. We validate our approach using two case studies of rail disruptions in two French towns in the suburbs of the Paris region. Results demonstrate that our method recovers the accessibility loss during disruptions much more efficiently than conventional replacement buses, restoring more accessibility with less driving distance, even without adding buses, just by appropriately extending existing lines and reallocating the existing fleet of buses.
@@ 2 Related Work
Due to its inherent complexity, Urban Rail Transit (URT) disruption management has been largely studied in the literature. Shalaby et al. [11] reviewed recovery models and algorithms for real-time railway disruption management, which involves timetable adjustment, substitute bus scheduling, and crew rescheduling.
@@ 2.2 Work Related to Replacement Buses
When an entire rail line is disrupted, an approach commonly employed in practice [6] is to deploy replacement buses (also called dedicated shuttle fleets) to travel across all stops of the disrupted rail line [20, 21]. Previous literature mainly focused on minimizing transit operating time, passenger discomfort, and system cost. In the same sense, Wang et al. [13] created a model for bus routes optimization for overall time driving minimization, whereas Cadarso et al. [22] designed an optimization model of bus bridging schedules and vehicle rescheduling from a comprehensive point of view, considering minimization of time of recovery, passenger discomfort, and system costs. Luo and Xu [23] incorporated the stochastic nature of passenger demand and existing rail and bus routes’ backup capacities into a stochastic programming model to minimize expected unmet passenger requirements. To cope with uncertain and heterogeneous bus traveling times, Liang et al. [24] developed a passenger flow and operational cost optimization model.
2.3 Accessibility Gap in Existing URT Disruption Literature
Despite the amount and high quality of previous work on Urban Rail Transit disruption remediation, all previous studies have overlooked the most crucial factor of a transportation system, i.e., accessibility [7]. This aspect is particularly important in suburban regions, where accessibility often tightly depends on one or few rail lines. Due to the scarcity of public transport alternatives, a disruption of a rail line can severely impact the accessibility of suburbs, preventing residents from performing their daily activities.
While accessibility is a well-established concept in transportation planning, our investigation confirms that URT disruption management research has not explicitly incorporated it as an optimization objective. Liu et al. [8] define accessibility as “the potential of opportunities for interaction” and recognize that transit systems face high risk of degradation from external disruptions, affecting their ability to deliver reliable accessibility. Their work introduces measures to detect disruptions’ impacts, but focuses on measurement rather than intervention. Similarly, Nalin et al. [25] emphasize the necessity of assessing the actual performance of public transportation rather than relying on schedules when evaluating accessibility, yet do not propose recovery strategies. This critical gap in the literature, the absence of optimization methods to restore accessibility through network redesign, constitutes our paper’s core contribution.
@@ 2.4 Our Contribution
For this reason, to the best of our knowledge, we are the first to propose a bus network redesign method with the explicit objective of optimizing accessibility restoration during Urban Rail Transit (URT) disruptions, particularly suited for suburbs with scarce public transport resources.
@@ 3.1 Tessellation of the Study Area
We partition the study area using a regular tessellation, where the geometric center of each tile is defined as a centroid. Let $\pazocal{C}$ denote the set of all centroids in the study area. A traveler can choose from various transportation modes for a trip between any pair of centroids. For instance, a traveler may walk the entire distance at speed $v_{walking}$, or they could walk from the origin centroid $c_{i}$ to a PT station, then take PT to another station, and finally walk from there to the destination centroid $c_{j}$. Observe that while actual trips may be longer or shorter than centroid-to-centroid distances, these variations tend to balance out across the PT network. This is why we approximate all travel as centroid-to-centroid.
@@ 3.2 Graph Model of PT Network
We model the original PT structure before disruption as graph $\pazocal{G}^{o}=(\pazocal{V}^{o},\pazocal{E}^{o})$. $\pazocal{G}^{o}$ is composed of multiple PT lines. Each line $\ell$ has a headway $t_{\ell}$, defined as the distance between two consecutive vehicles expressed in time [26]. The headway can be calculated using the formula $t_{\ell}=\frac{T_{\ell}}{N_{\ell}}$, where $T_{\ell}$ represents the round trip time along the entire line, and $N_{\ell}$ denotes the number of vehicles operating on that line (Eq. 4.4 of [26]). A line consists of a sequence of stations connected by edges. The average waiting time for line $\ell$ is $\frac{t_{\ell}}{2}$ [27], We define $t(s_{j},s_{j+1})$, the time taken to travel between any two successive stations $s_{j}$ and $s_{j+1}$, and $t_{s_{j}}$, the dwell time at station $s_{j}$. For interchange at station $s_{j}$ from line $\ell$ to line $\ell^{\prime}$, we consider an average waiting time $\frac{t_{\ell^{\prime}}}{2}$ at station $s_{j}$.
@@ 3.3 Problem Statement
In the face of urban rail system disruptions, our research aims to maximize accessibility through a redesign of bus lines. We extend operating bus routes allowing buses of one line to serve both stops of the original line and consolidation points at which passengers concerned by the disruption can be picked up and dropped off. Our optimization problem makes the following five decisions: (a) To which candidate consolidation points should passengers from the disrupted rail stations be suggested to go? (b) To which bus lines should the consolidation nodes be assigned? (c) Which bus line should be extended in order to serve such passengers? (d) When a bus serves multiple consolidation points, what is the optimal order of visits? (e) How to re-assign the existing fleet of buses and allocate additional vehicles (if available) among the original and extended lines?
The practical implementation of this approach assumes that drivers can adapt to revised routes through real-time navigation systems commonly employed by transit operators worldwide. These systems, such as onboard consoles or tablets, provide turn-by-turn navigation and can rapidly deploy route changes during disruptions with minimal additional hardware requirements.
@@
As illustrated in Fig. 2, for any given line $\ell$, let $t_{\ell}^{a}$ and $t_{\ell}^{b}$ represent its terminals. We can activate an extended version of $\ell$ from terminal $t_{\ell}^{b}$, which will serve the stops of the regular $\ell$ as well as consolidation nodes that will be associated with this line. Let’s denote this extended line as $\ell^{EX,b}$ (red line). Similarly, we can define $\ell^{EX,a}$ (blue line). Let $\pazocal{L}_{\text{REG}}=\{\ell_{1},\ell_{2},$$\ell_{\ell}\}$ be the set of regular lines. Let $\pazocal{L}_{\text{EX}}=\{\ell^{EX,a},\ell^{EX,b}|\ell\in\pazocal{L}_{\text{REG}}\}$ the set of all potential extended bus lines.
Table 1: Mathematical notation for the optimization model.
@@
$n$ Consolidation node index.
$d$ Demand point index.
$i,j$ Indices for nodes.
$\ell$ Bus line index.
@@
$\pazocal{N}$ Set of consolidation nodes, i.e., boarding locations for affected passengers.
$\pazocal{D}$ Set of disrupted rail stations, where affected passengers are gathered.
$\pazocal{C}$ Set of centroids in the study area.
$\pazocal{L}=\pazocal{L}_{\text{REG}}\cup\pazocal{L}_{\text{EX}}$ Set of all bus lines in the network, where $\pazocal{L}_{\text{REG}}$ are regular lines and $\pazocal{L}_{\text{EX}}$ are extended lines proposed by the model.
@@
$q_{d}$ Passenger demand at disrupted rail station $d$, $\forall d\in\pazocal{D}$.
capcap Passenger capacity per bus.
$N^{o}_{\ell}$ Number of buses on regular line $\ell$, $\forall\ell\in\pazocal{L}_{\text{REG}}$.
@@
$D_{max}$ Maximum distance allowed from disrupted station to consolidation node.
@@
$\pazocal{G}$ PT structure.
$n_{\ell}\in\mathbb{Z}^{+}$ Additional vehicles per unit time on line $\ell$ or its extensions, $\forall\ell\in\pazocal{L}_{\text{REG}}$.
$x_{\ell}\in\mathbb{Z}^{+}$ Vehicles per unit time on line $\ell$, $\forall\ell\in\pazocal{L}$.
$y_{\ell^{EX}}\in\{0,1\}$ 1 if extended line $\ell^{EX}\in\pazocal{L}_{\text{EX}}$ is activated, 0 otherwise.
$w_{n\ell}\in\{0,1\}$ 1 if consolidation point $n$ is assigned to extended line $\ell$, 0 otherwise.
$z_{dn}\in\{0,1\}$ 1 if disrupted station $d$ is assigned to consolidation $n$, 0 otherwise.
$u_{i}^{\ell}\in\mathbb{Z}^{+}$ Positional order of node $i$ in line $\ell$ (0 if not visited).
$X_{ij}^{\ell}\in\{0,1\}$ 1 if line $\ell$ connects $i$ and $j$, 0 otherwise.
@@ 3.4.1 Objective function
where $O_{c_{j}}$ is the number of opportunities within the tile with centroid $c_{j}$ and $T_{\pazocal{G}}(c_{i},c_{j})$ is the shortest travel time between centroids $c_{i}$ and $c_{j}$, considering that trips can combine multiple lines and can also include walk from $c_{i}$ to $c_{j}$ or to/from some intermediate stops; Subscript $\pazocal{G}$ indicates that travel time depends on PT structure $\pazocal{G}$. To implement this accessibility measure, we partition the study area using a regular tessellation with 1-kilometer grid cells, where centroids are defined as geometric centers of each tile. The 1-kilometer grid size balances computational efficiency with spatial precision, representing a reasonable walking distance while capturing accessibility variations in suburban areas with sparse public transport networks. Note that $O_{c_{j}}$ is an input parameter, exogenously determined. Indeed, if the PT structure changes (for instance some edges are no more available because of disruption), some shortest paths may change accordingly, resulting in a change in travel time $TG(c_{i},c_{j})$. Let $\pazocal{G}^{o}$ be the original PT structure before disruption, and $\pazocal{G}^{\text{DISR}}$ be the disrupted PT structure. Then, it’s easy to show that:
@@ 3.4.2 Constraints
$\sum_{\ell\in\pazocal{L}_{\text{EX}}}w_{n\ell}\leq 1,$ $\forall n\in\pazocal{N}$ (4)
@@
$z_{dn}\leq\sum_{\ell\in\pazocal{L}_{\text{EX}}}w_{n\ell},$ $\forall n\in\pazocal{N},\forall d\in\pazocal{D}$ (6)
@@
Constraints Eq. 4 to Eq. 7 handle the allocation relationship between disrupted rail stations and consolidation nodes: Constraint Eq. 4 ensures each consolidation node is assigned to at most one extended line; Constraint Eq. 5 guarantees each disrupted rail station is allocated to exactly one consolidation node; Constraint Eq. 6 ensures disrupted rail stations are only assigned to stops served by extended lines; Constraint Eq. 7 limits the distance between a disrupted rail station and its assigned consolidation node to within the maximum allowable distance.
@@
$\begin{aligned} &\sum_{j\in\pazocal{N}}X_{t_{\ell}^{a}j}^{\ell^{EX,a}}=y_{\ell^{EX,a}};\\ &\sum_{j\in\pazocal{N}}X_{t_{\ell}^{b}j}^{\ell^{EX,b}}=y_{\ell^{EX,b}}\end{aligned}$ $\forall\ell\in\pazocal{L}_{\text{REG}}$ (8)
$\sum_{j\in\pazocal{N}}X_{ij}^{\ell}=\sum_{j\in\pazocal{N}}X_{ji}^{l},$ $\forall i\in\pazocal{N},\forall\ell\in\pazocal{L}_{\text{EX}}$ (9)
$\sum_{j\in\pazocal{N}}X_{ij}^{\ell}\leq w_{i\ell},$ $\forall i\in\pazocal{N},\forall\ell\in\pazocal{L}_{\text{EX}}$ (10)
$\sum_{i\in\pazocal{N}}X_{ij}^{\ell}\leq w_{j\ell},$ $\forall j\in\pazocal{N},\forall\ell\in\pazocal{L}_{\text{EX}}$ (11)
$\begin{aligned} &u_{i}^{\ell}-u_{j}^{\ell}+(|\pazocal{N}|+2)\\ &\quad\cdot X_{ij}^{\ell}\leq(|\pazocal{N}|+1)\cdot w_{i\ell}\end{aligned},$ $\begin{aligned} &\forall i,j\in\pazocal{N},i\neq j,\\ &\forall\ell\in\pazocal{L}_{\text{EX}}\end{aligned}$ (12)
$2\cdot w_{i\ell}\leq u_{i}^{\ell}\leq(|\pazocal{N}|+1)\cdot w_{i\ell},$ $\forall i\in\pazocal{N},\forall\ell\in\pazocal{L}_{\text{EX}}$ (13)
$u_{t_{\ell}^{a}}^{\ell^{EX,a}}=y_{\ell^{EX,a}}~;u_{t_{\ell}^{b}}^{\ell^{EX,b}}=y_{\ell^{EX,b}}$ $\forall\ell\in\pazocal{L}_{\text{REG}}$ (14)
Constraints Eq. 8 to Eq. 14 design the line extensions: Constraint Eq. 8 defines the starting point constraints for line extensions. As shown in Figure 2, each regular line $\ell$ has two terminals $t_{\ell}^{a}$ and $t_{\ell}^{b}$. The first term $\sum_{j\in\pazocal{N}}X_{t_{\ell}^{a}j}^{\ell^{EX,a}}=y_{\ell^{EX,a}}$ ensures that if line extension from terminal $t_{\ell}^{a}$ is activated, there must exist exactly one direct connection from $t_{\ell}^{a}$ to a consolidation node. Similarly, the second term $\sum_{j\in\pazocal{N}}X_{t_{\ell}^{b}j}^{\ell^{EX,b}}=y_{\ell^{EX,b}}$ addresses the extension from terminal $t_{\ell}^{b}$. Constraint Eq. 9 maintains flow conservation at consolidation nodes and ensures that the number of incoming and outgoing connections is equal. Constraints Eq. 10 and Eq. 11 limit each node to have at most one outgoing and one incoming edge in each line, respectively, preserving the linear structure of the line. Constraint Eq. 12 implements the Miller-Tucker-Zemlin subtour elimination constraint [28], ensuring that the line forms a single path without cycles among consolidation nodes. Lastly, constraint Eq. 13 sets the bounds for node position variables, maintaining sequential order, while Constraint Eq. 14 specifically designates the starting terminal.
@@
$x_{\ell^{EX,a}}\geq y_{\ell^{EX,a}}~;x_{\ell^{EX,b}}\geq y_{\ell^{EX,b}}$ $\forall\ell\in\pazocal{L}_{\text{REG}}$ (15)
$x_{\ell^{EX,a}}\leq M\cdot y_{\ell^{EX,a}}~;x_{\ell^{EX,b}}\leq M\cdot y_{\ell^{EX,b}}$ $\forall\ell\in\pazocal{L}_{\text{REG}}$ (16)
$cap\cdot x_{\ell}\geq\sum_{d\in\pazocal{D}}\sum_{n\in\pazocal{N}}q_{d}\cdot z_{dn}\cdot w_{n\ell},$ $\forall\ell\in\pazocal{L}_{\text{EX}}$ (17)
$n_{\ell}=x_{\ell}+x_{\ell^{EX,a}}+x_{\ell^{EX,b}}-N^{o}_{\ell},$ $\forall\ell\in\pazocal{L}_{\text{REG}}$ (18)
$\sum_{\ell\in\pazocal{L}}n_{\ell}\leq N_{max}$ (19)
Constraints Eq. 15 to Eq. 19 deal with vehicle allocation and line selection: Constraint Eq. 15 and Constraint Eq. 16 ensure vehicles are only assigned to a line extension if this extended line is activated; Constraint Eq. 17 guarantees the vehicle capacity of each extended line meets demand; Note that $x_{\ell}$ indicates the number of vehicles operating on the extended line $\ell$ and $z_{dn}\cdot w_{n\ell}$ is one if and only if disrupted rail station $d$ is served by extended line $\ell$; $q_{d}$ can be estimated through analysis of historical passenger flow data and real-time monitoring etc. In our numerical experiments, we employed statistical models; this constraint makes the problem non-linear. Constraint Eq. 18 governs the allocation of vehicles across the transit network. For each regular line $\ell$, it ensures that the number of vehicles on the regular line ($x_{\ell}$) equals the initial fleet size ($N^{o}_{\ell}$) plus additional vehicles fleet size ($n_{\ell}$), minus the number of vehicles on extended lines. Constraint Eq. 19 expresses a budget for additional buses to be added. Note that the formulation allows us to reduce the number of buses operating on regular lines in our solution with respect to the PT design before disruption. It may indeed be useful to slightly reduce the service on regular lines to compensate for the loss of accessibility due to disruption.